# Can AI Prompt Humans? Multimodal Agents Prompt Players’ Game Actions and Show Consequences to Raise Sustainability Awareness

### v1

@@ 3.2.2. Intent to Action: Player Intent Detection Mechanism
Figure 7. The AI-driven dialogue system achieves intelligent, dynamic NPC-player interactions through multi-stage processing. Initially, the AI-Driven NPC initiates dialogue based on pre-defined NPC character profiles. After player input, an intent detection agent analyzes user intentions, incorporating conversation context stored in a database. The system then evaluates the detected intent in a core decision phase, selecting the most appropriate response strategy based on specific circumstances: (1 Generating a new response, (2) Returning a predefined answer, or (3) Providing a specific in-game item
@@ 3.3. Technical Architecture of EcoEcho
Figure 8. Integrated Technical Architecture of EcoEcho: Combining React Frontend, Express Backend, AI-Driven Prompt Engineering and Dialogue Generation using Llama 70B, and Specialized AI Tools (ElevenLabs for Voiceovers, Suno for Background Music, DALL-E for Visuals) for Immersive Gaming Experience
@@ 4.1. Participants
Table 1. Participant Demographics by Age, Gender, Education, Occupation, and Geographical Distribution(N=23)
ID Age Gender Education Occupation Location
1 18-24 Male Undergraduate Student USA
2 25-34 Female Postgraduate Student Canada
3 25-34 Female Postgraduate Student USA
4 25-34 Female Postgraduate Student USA
5 25-34 Female Undergraduate Employee China
6 18-24 Female Undergraduate Employee China
7 25-34 Female Undergraduate Student China
8 18-24 Female Undergraduate Other China
9 18-24 Female Undergraduate Student China
10 25-34 Male Postgraduate Other China
11 25-34 Female Postgraduate Other New Zealand
12 25-34 Female Postgraduate Student New Zealand
13 25-34 Male Postgraduate Student New Zealand
14 25-34 Male Postgraduate Student New Zealand
15 25-34 Female Postgraduate Student China
16 25-34 Female Postgraduate Student New Zealand
17 18-24 Male Undergraduate Student New Zealand
18 25-34 Female Postgraduate Student New Zealand
19 18-24 Female Postgraduate Student China
20 18-24 Male Undergraduate Student China
21 18-24 Female Undergraduate Student USA
22 18-24 Female Postgraduate Student USA
23 18-24 Male Undergraduate Student China
@@
Figure 9. Study Procedure: Pre-Survey (2 Weeks Before Playing): Collect demographic information, awareness of environmental sustainability, and AI usage in daily life. Game Session (Playing the Game): Track players’ attitudes and behaviors toward sustainability within the game. Post-Survey (After Playing): Evaluate changes in players’ sustainability attitudes and perceptions through questionnaires and semi-structured interviews.
@@ 4.2. User Study Procedure
A pre-survey was conducted two weeks before the gaming session online to explore participants’ backgrounds and environmental sustainability awareness. In the pre-survey session, we used the scales of the General Ecological Behavior (Kaiser and Wilson, 2000) and New Ecological Paradigm (Dunlap et al., 2000), with a five-point scale and adapting them to focus on energy issues within the context of sustainability as related to our game. We also assessed participants’ use of generative AI in their daily lives, as generative AI was used in the game’s art and music, and the NPCs were powered by large language models.
@@
In the post-survey session, we used a questionnaire, similar to the pre-survey, to assess participants’ environmental sustainability awareness. We also explored their satisfaction with various aspects of the game to evaluate whether their attitudes toward sustainability had changed after playing and whether the use of generative AI in the game had any impact. Following this, we conducted semi-structured interviews to qualitatively explore how the AI agents and the game itself influenced their attitudes toward energy issues throughout the gameplay.
@@ 5.1.2. Interacting with NPCs
As players make more unsustainable choices, the game’s environment and scenarios begin to shift. The Agent employs a feedback mechanism in which players’ decisions directly impact the game world’s environment. For example, the scenery outside the window changes from lush, green trees to smoke-belching factories (see Figure 5). At this point, players begin to realize the negative consequences of their early decisions and start to reflect on their actions.
@@ 5.2. Sustainable Awareness and Behavior Before and After the Game
Table 2. Shapiro-Wilk test results for NEP and GEB data.
(a) (a) General Ecological Behavior Pre vs. Post
(b) (b) New Ecological Paradigm Pre vs. Post
(c) (c) Changes in the GEB
(d) (d) Changes in the NEP
Figure 10. Impact of the gamified prompt intervention on ecological intended behavior and attitudes: (a) shows that post-intervention GEB scores demonstrate a significant increase in median values, indicating an overall improvement in participants’ ecological intended behaviors. The reduced interquartile range suggests more consistent intended behavioral changes across participants. (b) indicates that NEP scores post-intervention exhibit a slight shift towards higher values, reflecting a modest but positive change in environmental attitudes. However, the variability in scores remains similar, indicating diverse responses to the intervention. (c) highlights that the majority of participants experienced an increase in GEB scores post-intervention, as evidenced by the higher count of positive ranks, suggesting that the intervention effectively enhanced participants’ ecological intended behaviors, with fewer participants showing no change or a decrease. (d) shows that while there is a positive shift in NEP scores, the number of participants with increased ranks is slightly lower compared to GEB, indicating that while the intervention influenced environmental attitudes, the effect was less pronounced compared to intended behavioral changes.
5.2.1. New Ecological Paradigm
@@ 5.3.2. Post-Game Feedback
Figure 11. Heatmap representing the voting patterns of 23 players across four rounds of the gamified assessment. Each cell in the heatmap shows the Likert scale score (ranging from 0 to 5) that players assigned at each round, with ”First time” to ”Fourth time” indicating the sequential rounds of voting. The color intensity reflects the value of the scores, where darker shades correspond to higher scores. The y-axis labels represent the 23 individual players, numbered from 1 to 23. This visualization highlights how player attitudes evolved over time, with noticeable shifts in voting action.
@@ 5.4. Qualitative Feedback On Players’ Perception After The Gameplay
Table 3. Questions on Game Experience and Sustainable Awareness and Behavior
@@ 6.2. Action Before Attitude: Games Drive Behavior Change Ahead of Mindset Shift
In EcoEcho, we adopted a narrative approach opposite to that of traditional persuasive games, where good deeds lead to positive outcomes. Instead, participants must take actions in their character’s interest, which ultimately results in negative consequences. In fact, many video games incorporate this approach by presenting players with moral dilemmas, such as in Detroit: Become Human, Spec Ops: The Line, BioShock, and This War of Mine. Jørgensen  (Jørgensen, 2016) studied the emotional impact on participants after playing the third-person military shooter game Spec Ops: The Line, coining the term ”The Positive Discomfort.” This emotion has the potential to raise awareness by provoking participants to reflect and question their culturally ingrained and fundamental values. In the interviews, participants mentioned experiencing negative emotions such as guilt (P1, P2, P7, P5, P15, P21) and fear (P2, P3, P4, P8, P14, P20) as a result of their actions in the game. Additionally, they were able to connect their in-game actions with real-world consequences. As P9 noted, ”I felt that just like in daily life, if we don’t prioritize clean energy, our world could one day end up like the (game’s) outcome.” Interestingly, we also found that some participants were not affected by the negative outcomes depicted in the game at all, as P12 stated: ”It wasn’t me who caused all this, it was my character.” In fact, this mindset is widely observed in role-playing games and is known as ”alibi,” (Bowman, 2021) where players distance themselves from their in-game actions by attributing them to their character. Alibi shields some players from the long-term impact of negative emotions after the game ends, but it also diminishes the likelihood of engaging in reflective thinking.
However, in our analysis of the pre- and post-test results, we found that EcoEcho had varying degrees of impact on participants’ intended behavior and attitudes toward sustainability. There was a correlation between participants’ environmental attitudes and sustainable behaviors before playing the game—where positive environmental attitudes tended to promote sustainable behaviors, which aligns with prior studies (Kesenheimer and Greitemeyer, 2021; Balaskas et al., 2023)—this relationship did not result in parallel changes after playing EcoEcho. The post-test results showed a significant increase in participants’ scores on the General Ecological Behavior scale, suggesting that they believed they would engage in more sustainable behaviors. However, their environmental attitudes did not show a significant change.
As some studies (Zhou et al., 2024; Hirsh, 2010) indicated, environmental attitudes are related to personality traits like ’Openness’ and ’Agreeableness,’ as well as a preference for democratic values. These traits are long-standing (Biesanz et al., 2003) and difficult to change dramatically in a short period of time. On a positive note, through the game, we were able to prompt players’ intended sustainable behaviors, with more participants expressing that they would be more mindful of sustainability in their daily lives even though they appear to retain their core personalities. However, as Verplanken et al.’s research (Verplanken and Orbell, 2022) pointed out, long-term behavior changes have the potential to solidify into habits, and these habits, in turn, can influence personal attitudes. In other words, if players are prompted to engage in sustainable behavior over the long term, their environmental attitudes may also undergo positive changes. Besides, in the pre-survey, we found that our participants generally already had a strong sense of sustainability, suggesting that the younger generation inherently possesses a heightened environmental awareness, which also aligns with previous research findings (Adu et al., 2021; Lazăr et al., 2022).

### v4

@@ 3.2.2. Intent to Action: Player Intent Detection Mechanism
Figure 6. The AI-driven dialogue system achieves intelligent, dynamic NPC-player interactions through multi-stage processing. Initially, the AI-Driven NPC initiates dialogue based on pre-defined NPC character profiles. After player input, an intent detection agent analyzes user intentions, incorporating conversation context stored in a database. The system then evaluates the detected intent in a core decision phase, selecting the most appropriate response strategy based on specific circumstances: (1 Generating a new response, (2) Returning a predefined answer, or (3) Providing a specific in-game item.
@@ 3.3. Technical Architecture of EcoEcho
Figure 7. System architecture of EcoEcho, an AI-enhanced educational game platform. The system integrates frontend development (React), backend services (Express), and multiple AI components: Llama 70B for NPC dialogue generation, along with supplementary AI tools for audio-visual content creation, connected through API interfaces with human-in-the-loop feedback mechanisms.
@@ 4.1. Participants Demographic
2 ID Age Gender Education Occupation Location 1 18-24 Male U Student USA 2 25-34 Female G Student Canada 3 25-34 Female G Student USA 4 25-34 Female G Student USA 5 25-34 Female U Employee China 6 18-24 Female U Employee China 7 25-34 Female U Student China 8 18-24 Female U Other China 9 18-24 Female U Student China 10 25-34 Male G Other China 11 25-34 Female G Other New Zealand 12 25-34 Female G Student New Zealand 13 25-34 Male G Student New Zealand 14 25-34 Male G Student New Zealand 15 25-34 Female G Student China 16 25-34 Female G Student New Zealand 17 18-24 Male U Student New Zealand 18 25-34 Female G Student New Zealand 19 18-24 Female G Student China 20 18-24 Male U Student China 21 18-24 Female U Student USA 22 18-24 Female G Student USA 23 18-24 Male U Student China
Table 1. Participant Demographics by Age, Gender (U - Undergraduate; G - Postgraduate), Education, Occupation, and Geographical Distribution (N=23) in the Pre-Survey.
@@ 4.2. User Study Procedure
A pre-survey was conducted two weeks before the gaming session to explore participants’ backgrounds and their awareness of environmental sustainability. The General Ecological Behavior (GEB) scale was used to assess participants’ pro-environmental actions before playing the game, such as energy-saving habits and waste reduction, while the New Ecological Paradigm (NEP) scale evaluated their environmental attitudes, including beliefs about ecological limits and human-nature interactions. Both scales were adapted to focus on energy issues relevant to the game. Additionally, we assessed participants’ use of generative AI in daily life, given its role in the game’s art, music, and NPCs powered by large language models.
Figure 8. Timeline of the three-phase study methodology examining players’ environmental sustainability awareness. The study consists of a pre-survey phase assessing participants’ baseline characteristics and knowledge, followed by a gameplay phase monitoring sustainability-related behaviors, and concluding with a post-gameplay evaluation combining questionnaire and interview data to measure attitudinal changes.
@@
In the post-survey session, we used a questionnaire, similar to the pre-survey, to assess participants’ environmental sustainability awareness, including intended pro-environmental actions and environmental attitudes after playing the game. We also explored their satisfaction with various aspects of the game to evaluate whether their attitudes toward sustainability had changed after playing and whether the use of generative AI in the game had any impact. Following this, we conducted semi-structured interviews to qualitatively explore how the AI agents and the game itself influenced their attitudes toward energy issues throughout the gameplay.
4.3. Analysis Protocol
4.3.1. Quantitative Analysis Procedure
To analyze the impact of the game on participants’ environmental attitudes and behaviors, we used two established scales: NEP and GEB scales. The NEP scale evaluated their environmental attitudes and the GEB scale assessed their intended sustainable behaviors. For the NEP scale, the Shapiro-Wilk test confirmed that both pre-test and post-test scores were approximately normally distributed. Thus, a paired samples t-test was conducted to compare pre- and post-test scores. For the GEB scale, while pre-test scores were normally distributed, post-test scores were not. Consequently, the Wilcoxon signed-rank test was used to examine changes in GEB scores.
We also evaluated participants’ attitudes toward sustainable energy through in-game assessments, conducted via voting on whether to support sustainable energy development at four stages of the game. Participants were given five votes during each stage and informed that insufficient supporting votes would result in the failure of the vote, reflecting their stance on sustainable energy policies. Descriptive analysis was then used to observe how their voting behaviors changed across these stages.
4.3.2. Qualitative Analysis Procedure
The qualitative analysis was conducted using Thematic Analysis (TA), following a structured, collaborative approach between two researchers. First, we independently read through all interview transcripts to familiarize ourselves with the content and context. During this phase, we noted interesting snippets and identified recurring patterns. This familiarization process took approximately two days.
Next, we employed an open coding method to break down the transcripts into smaller text segments. Descriptive labels (codes) were assigned to each segment to capture key information relevant to the study’s focus, such as climate attitudes or gameplay experiences. For example, in Table 3, responses to Question (1) might be coded as ”emotional resonance with NPC dialogues,” while responses to Question (9) could be coded as ”change in behavioral intent” or ”real-life sustainable actions.” During this stage, we collaboratively developed an initial set of codes, iteratively refining their definitions until a mutual agreement was reached.
Once a consistent coding framework was established, each of us independently applied the codes to all textual data. We continued to refine and merge existing codes, grouping similar or related codes into broader themes. Any overlapping or unclear themes were adjusted through discussion, and this iterative process continued until no new categories could be found. Finally, one of us defined clear boundaries and meanings for each theme and assigned concise, representative names. These themes were supported by direct quotes from participants to ensure transparency and rigor in the qualitative analysis.
@@ 5.1.2. Interacting with NPCs
As players make more unsustainable choices, the game’s environment and scenarios begin to shift. The Agent employs a feedback mechanism in which players’ decisions directly impact the game world’s environment. For example, the scenery outside the window changes from lush, green trees to smoke-belching factories (see Figure LABEL:fig:fig4). At this point, players begin to realize the negative consequences of their early decisions and start to reflect on their actions.
@@ 5.2. Sustainable Awareness and Behavior Before and After the Game
(a) (a) Changes in General Ecological Behavior (GEB) and New Ecological Paradigm (NEP) Pre- and Post- Scores.
(b) (b) Distribution of Positive, Negative, and No Change Ranks in GEB and NEP Scores.
Figure 9. Impact of the gamified prompt intervention on ecological intended behavior and attitudes:(a) demonstrates that post-intervention GEB and NEP scores exhibit noticeable changes. The GEB scores show a significant increase in median values with a reduced interquartile range, indicating an overall improvement and more consistent changes in participants’ ecological intended behaviors. The NEP scores, while showing a modest shift toward higher values, reflect a positive change in environmental attitudes. However, the variability in NEP scores remains similar, suggesting diverse participant responses to the intervention. (b) highlights the rank changes post-intervention, showing that the majority of participants experienced an increase in GEB scores, evidenced by the higher count of positive ranks, indicating the intervention effectively enhanced participants’ ecological intended behaviors. For NEP scores, while there is a positive shift, the number of participants with increased ranks is slightly lower compared to GEB, suggesting that the intervention’s impact on environmental attitudes, though positive, was less pronounced compared to intended behavioral changes.
5.2.1. New Ecological Paradigm
@@
Table 2. A Statistical Examination of Environmental Metrics: Shapiro-Wilk Test Results (W=0.977-0.894) Demonstrating Normal Distribution in NEP (Pre/Post) and GEB Pre-test with Non-normal Distribution in GEB Post-test Data
@@ 5.3.1. In-Game Voting Results
Figure 10. Heatmap representing the voting patterns of 23 players across four rounds of the gamified assessment. Each cell in the heatmap shows the Likert scale score (ranging from 0 to 5) that players assigned at each round, with ”First time” to ”Fourth time” indicating the sequential rounds of voting. The color intensity reflects the value of the scores, where darker shades correspond to higher scores. The x-axis labels represent the 23 individual players, numbered from 1 to 23. This visualization highlights how player attitudes evolved over time, with noticeable shifts in voting action.
@@ 5.4. Qualitative Feedback On Players’ Perception After The Gameplay
Table 3. Semi-structured Interview Questions Assessing Gaming Experience and Environmental Sustainability Outcomes
@@ 6.2. Action Before Attitude: Games Drive Behavior Change Ahead of Mindset Shift
In EcoEcho, we adopted a narrative approach opposite to that of traditional persuasive games, where good deeds lead to positive outcomes. Instead, participants must take actions in their character’s interest, which ultimately results in negative consequences. Inspired by the ”Exposition” from the RECIPE framework (Nicholson, 2015), EcoEcho visualizes the consequences of actions in an easily perceivable way. By showcasing the outcomes of players’ “Choices”, this approach makes abstract sustainability challenges, such as climate change, more concrete, fostering players’ “Reflection” and potentially encouraging behavior change. Many video games incorporated and modified this approach by presenting players with moral dilemmas and bad consequences, such as in Detroit: Become Human, Spec Ops: The Line, BioShock, and This War of Mine. Jørgensen  (Jørgensen, 2016) studied the emotional impact on participants after playing the third-person military shooter game Spec Ops: The Line, coining the term ”The Positive Discomfort.” This emotion has the potential to raise awareness by provoking participants to reflect and question their culturally ingrained and fundamental values. In the interviews, participants mentioned experiencing negative emotions such as guilt (P1, P2, P7, P5, P15, P21) and fear (P2, P3, P4, P8, P14, P20) as a result of their actions in the game. Additionally, they were able to connect their in-game actions with real-world consequences. As P9 noted, ”I felt that just like in daily life, if we don’t prioritize clean energy, our world could one day end up like the (game’s) outcome.” Interestingly, we also found that some participants were not affected by the negative outcomes depicted in the game at all, as P12 stated: ”It wasn’t me who caused all this, it was my character.” In fact, this mindset is widely observed in role-playing games and is known as ”alibi,” (Bowman, 2021) where players distance themselves from their in-game actions by attributing them to their character. Alibi shields some players from the long-term impact of negative emotions after the game ends, but it also diminishes the likelihood of engaging in reflective thinking.
Previous research shown the subtle relationships between sustainability awareness, attitude, and behavior. For example, Arshad et al. (Arshad et al., 2021) found that awareness and attitudes do not always correlate, while Gao (Firmanshah et al., 2023) and Safari et al. (Cogut et al., 2019) reported a positive correlation between sustainability awareness and behavior. Our results align with these findings, as the improvement in intended behavior suggests a link to increased awareness, indicating that our game effectively enhanced participants’ sustainability awareness. However, no significant changes were observed in their environmental attitudes, suggesting that while awareness and attitudes may not be directly correlated, their interaction could occur indirectly through other factors, such as behavior. This implies that changes in behavior might precede shifts in attitudes, reflecting a phenomenon known as cognitive dissonance reduction, where individuals adjust their attitudes to align with their actions over time (Festinger, 1957). Alternatively, it may suggest that environmental attitudes are more resistant to change and require prolonged engagement or deeper emotional connections to evolve.
This raises an important question: Can the intended sustainable behaviors observed in our study translate into long-term real-life actions? Prior research by Douglas and Brauer (Douglas and Brauer, 2021) indicates that engaging in pro-environmental behaviors within a game context can encourage individuals to maintain these behaviors after gameplay, fostering habit formation and gradually influencing attitudes. According to the SHIFT model (White et al., 2019), games can effectively address key components of behavior change, including social influence, habit formation, and the tangibility of actions. By providing a structured yet engaging environment for practicing pro-environmental behaviors, games offer a low-risk space to experiment and internalize these actions. Repeated behaviors performed in-game may translate into real-world habits, which, over time, can contribute to shifts in attitudes and long-term behavioral change. Our qualitative findings support this notion, as many participants expressed greater mindfulness toward sustainability in their daily lives after playing the game. Additionally, our quantitative results, which showed a significant self-reported increase in sustainable behaviors, provide further evidence that these behaviors may persist over time. Verplanken et al. (Verplanken and Orbell, 2022) also emphasized that long-term behavioral changes can eventually solidify into habits, which in turn have the potential to shape attitudes. Based on these findings and the evidence from both our qualitative and quantitative results, it is plausible that the behavioral shifts prompted by our game could evolve into sustained habits, which may later contribute to positive attitude changes toward sustainability.
Thus, while our study observed limited immediate changes in attitudes, the combination of increased intended behaviors and participants’ self-reported mindfulness suggests the possibility of long-term effects. Sustained engagement in these behaviors could act as a bridge to eventual shifts in attitudes, aligning with existing research on the habit-attitude relationship. Future longitudinal studies could provide further insight into this potential trend.
@@ 6.4. From Dialogue to Game Actions: Applying Multi-modal Agents to Simulate Real-World Behaviors
6.5. Ethical Considerations
6.5.1. Risk of Manipulation
Our results demonstrate an increase in sustainability awareness through the use of AI-driven prompts to influence players’ decisions and behaviors in a simulated environment. However, we acknowledge the concern that similar algorithms could be misused to promote malicious agendas, exploit user vulnerabilities, or reinforce harmful behaviors (Brundage et al., 2018).
To mitigate these risks, we have implemented a strategy to ensure secure and responsible use of the technology:
  • Strict Prompt Engineering and Usage Boundaries: We designed strict prompt engineering protocols to ensure that the AI-driven NPCs operate within clearly defined ethical boundaries. These prompts are regularly reviewed and tested to prevent outputs that may lead to unintended or harmful consequences while maintaining player autonomy.
  • Content Vetting and Multilingual Support: The system supports both English and Chinese versions through Google Translation APIs. To ensure consistency and prevent unintended biases, all translated content is carefully vetted by human reviewers, minimizing risks of misinterpretation or harmful responses across languages.
  • User Privacy and Informed Consent: Participants in our study are fully informed about how their data will be used, and we prioritize privacy by anonymizing responses and avoiding the collection of sensitive personal data. These measures align with best practices for ethical research and data protection.
While these steps address immediate concerns, preventing misuse at scale requires a broader collaborative effort. We advocate for transparency in the development and deployment of AI technologies and suggest the following actions: Researchers should develop and implement robust prompt engineering practices to minimize the risk of harmful applications. Technology companies should take the lead in informing stakeholders about the risks and ethical considerations of AI technologies. Additionally, society, including independent organizations and regulatory bodies, should conduct external audits to ensure ethical and transparent AI design.
By combining technical safeguards with clear policies and collaborative oversight, we aim to ensure the secure and ethical application of this technology, preventing its misuse while maximizing its potential for positive societal impact.
6.5.2. The Unsustainable Side of GenAI
Another important question that requires serious consideration is the environmental impact of generative AI systems, particularly large language models, which consume significant energy  (Patterson et al., 2021). To align with the sustainability values promoted by our game, we made deliberate design choices to minimize resource usage. We use pre-trained, open-source models (e.g., LLaMA 70B), avoiding any type pf additional training, and adopt smaller, efficient architectures to reduce computational costs while maintaining interactive capabilities for player engagement.
While AI-driven content generation offers potential advantages, such as reducing the overhead of traditional game updates by enabling targeted and dynamic adaptations, we acknowledge that these benefits do not negate the broader challenges of resource-intensive AI applications. It is crucial to critically evaluate and mitigate the environmental impact of deploying such technologies. We recommend that future research and development focus on optimizing the energy efficiency of generative AI systems by exploring lightweight architectures, adopting carbon-aware computing strategies, and leveraging renewable energy sources. Additionally, developers and organizations should establish clear guidelines for responsible usage, such as limiting redundant computations and encouraging shared infrastructure to reduce the carbon footprint of AI applications. By integrating these sustainable practices into both the design and deployment phases, we can ensure that generative AI systems align more closely with ethical and environmental priorities.
6.6. Design Implications
6.6.1. Serious Games
Traditional serious games are sometimes criticized for having complex mechanics and high learning curves, which can limit engagement from the general public (Van Oostendorp et al., 2014). In EcoEcho, we simplified the interaction by using natural language. Participants engaged with NPCs through chatbot-like conversations, completing tasks driven by the narrative context. We found that such interactions could promote intended sustainable behaviors and encourage players to consider issues from multiple perspectives. This study proposes a new design method that integrates LLMs into game mechanics to drive interactions between NPCs and players. This approach reduces participants’ mental load in learning the game mechanics and has the potential to attract a broader, non-gaming audience to engage with sustainable serious games.
Besides, we used generative AI to facilitate rapid game development for EcoEcho. In practical settings such as schools, teachers often lack the skills to develop games or access necessary game assets, making it difficult to implement serious games tailored to specific educational needs (Stanitsas et al., 2019). By leveraging generative AI, teachers could more easily generate required game assets, create simple game code, and develop customized games at low cost, reducing barriers to adopting serious games in real-world applications. However, ethical concerns widely arise in this process, particularly regarding debates over copyright issues associated with AI-generated assets (Samuelson, 2023) and potential biases in AI-generated results (Srinivasan and Uchino, 2021). Therefore, we encourage the game development community to engage in discussions on how AI-generated materials can be ethically and effectively utilized in serious games to promote public well-being, and to work towards establishing relevant guidelines.
6.6.2. Non-Game Interventions
Our research demonstrates that AI-driven interventions in games can effectively influence intended behaviors and foster sustainability awareness. These insights can also be generalized to non-game contexts to maximize their impact in promoting sustainability.
Narrative Applications Outside of video games, EcoEcho’s narrative-driven approach can inspire the use of storytelling in other media to engage audiences on sustainability issues. For example, AI-powered digital platforms can adopt modular, narrative-driven designs similar to EcoEcho’s rich NPC interactions to personalize sustainability education. These platforms can foster a nuanced understanding of environmental challenges by adapting game-like storytelling techniques to contexts like e-learning or public awareness campaigns.
Policy Implications EcoEcho’s multimodal conversational AI systems demonstrate how AI can guide users toward sustainable behaviors through personalized, interactive experiences. Policymakers and organizations can harness these insights to develop AI-driven platforms for public awareness campaigns. For instance, a multimodal conversational AI system could assist users in calculating their carbon footprints through conversational prompts combined with visual feedback, offering actionable strategies to reduce emissions. Decision-support tools, modeled after EcoEcho’s dynamic response systems, can aid policymakers and businesses in evaluating the outcomes of sustainability initiatives (Cortès et al., 2000), such as adopting renewable energy or minimizing packaging waste. By analyzing user inputs and providing scenario-based feedback, these tools can support informed decision-making and environmental protection.
Use of GenAI GenAI systems can replicate EcoEcho’s multimodal approach to enhance user engagement and promote eco-friendly practices. By combining text, voice, and visual elements, AI-driven platforms can simulate real-world consequences of unsustainable actions, helping users visualize the impact of behaviors like excessive waste or high energy consumption. Additionally, GenAI could facilitate discussions among agents representing diverse perspectives, such as policymakers, community leaders, and scientists. These dialogues could present trade-offs, highlight diverse viewpoints and encourage participants to engage in collaborative problem-solving to identify feasible and balanced solutions to sustainability challenges.